OpenStudy (anonymous):

Let (p,q) be a point on the curve sqrt(x) + sqrt(y)=1 and dy/dx=-sqrt(y/x). Show that the tangent line to this curve at (p,q) has equation (x)sqrt(q) + (y)sqrt(p) = sqrt (pq). I can't seem to get the correct answer for the RHS.

OpenStudy (anonymous):

This is what I keep getting: \[x \sqrt{q} + y \sqrt{p} = p \sqrt{q} + q \sqrt{p}\]

ganeshie8 (ganeshie8):

Point : \(\large(p, q)\) Slope : \(\large -\frac{q}{p}\) Line : \(\large y - q = -\sqrt{\frac{q}{p}}(x-p)\)

OpenStudy (anonymous):

yep, I did that

ganeshie8 (ganeshie8):

simplifying gives : \(\large x\sqrt{q} + y \sqrt{p} = q\sqrt{p} + \sqrt{pq}\)

OpenStudy (anonymous):

the answer requires only \[\sqrt{pq}\] :/

ganeshie8 (ganeshie8):

yeah, not sure how to kick off that extra \(\large q\sqrt{p}\) from right hand side :/

OpenStudy (anonymous):

damn.. thanks anyway :)

ganeshie8 (ganeshie8):

np :) its not wise to call the text book has a mistake... but.. im not getting any other clue ...

OpenStudy (anonymous):

it's actually an assignment :P

OpenStudy (anonymous):

Oh, I've got it. Take \[\sqrt{pq}\] out as a common factor so you're left with: \[\sqrt{pq}(\sqrt{p} + \sqrt{q})\] Sub (p,q) into the equation given to so that: \[\sqrt{p} + \sqrt{q} = 1\] And the RHS will equal: \[\sqrt{pq}\]

ganeshie8 (ganeshie8):

brilliant !! xD